### v2

@@ 4.1.2 Implementation Details
To ensure a fair comparison with prior work, we have adopted the U-Net architecture as the backbone for our segmentation model. Following previous work [37, 2], we employ the Dice Similarity Coefficient (DSC), the Jaccard Index, the 95th percentile Hausdorff Distance (95HD), and the Average Surface Distance (ASD) to evaluate the segmentation accuracy of our medical image analysis model. During the training process, each batch consists of an equal mix of 12 labeled and 12 unlabeled data samples. We initialized the learning rate at 0.01 and used an SGD optimizer. The model was trained for 300 epochs under a learning rate scheduler. The original images were resized to vary between 0.5 and 2.0 times their original dimensions, and subjected to cropping and flipping to obtain their weakly perturbed versions.
4.2 Semi-supervised image segmentation in ACDC
@@
In the ACDC dataset, we compared our experimental results with those of multiple models, consistently achieving superior results. In comparison to the previous state-of-the-art (SOTA), our method demonstrated improvements across four metrics: DICE, Jaccard, 95HD, and ASD.
Method
1 case
3 case
7 case
DICE $\uparrow$ / Jaccard $\uparrow$ / 95HD $\downarrow$ / ASD $\downarrow$
SupBaseline 28.50/16.62/ - / - 47.83/37.01/31.16/12.62 79.41/68.11/9.35/2.70
UA-MT (MICCAI 2019) - / - / - / - 46.04/35.97 /20.08/7.75 81.65/70.64/6.88/2.02
DTC (AAAI 2021) - / - / - / - 56.90/45.67/ 23.36/7.39 84.29/73.92/12.81/4.01
SS-Net(MICCAI 2022) - / - / - / - 65.82/55.38/6.67/2.28 86.78/77.67/6.07/1.40
BCP (CVPR 2023) - / - / - / - 87.59/78.67/1.90/0.67 88.84/80.62/ 3.98/1.17
Unimatch (CVPR 2023) 85.43/74.57/ - / - 88.86/ 79.95 / - / - 89.85/ 81.57 / - / -
CrossMatch(Arxiv 2024) - / - / - / - 88.27/80.17/1.53/0.46 89.08/81.44/1.52/0.52
Ours 86.25$\pm$2.50/75.82/1.11/0.21 89.53$\pm$2.15/81.07/0.94 /0.17 90.40$\pm$1.68/82.48/0.86/0.14
Table 1: Comparison with seven different models on the ACDC dataset, using 1, 3, and 7 labeled cases. SupBaseline denotes the performance achieved by supervised leveraging solely labeled imagery within the scope of the study. Our method, AIGCMatch, has the best performance.
@@
It should be noted that, compared to previous methodologies, our approach has yielded outcomes equivalent or superior to those of other models that have used 10% labeled data, although it only employs 5% labeled data. Furthermore, compared to previous methods that labeled 5% labeled data, our approach has achieved comparable or superior outcomes with only 1% of the data, which has often been overlooked due to their poor performance in previous studies. This substantiates the superior performance and substantial potential of AIGCMatch in the domain of medical image segmentation.
4.3 Semi-supervised image segmentation in ISIC-2017
The ISIC-2017 dataset is widely recognized as challenging[43], In our research, we conducted a thorough performance comparison on the ISIC-2017 dataset, evaluating our method against a baseline model. As illustrated in Table 2, the Weak-to-Strong Consistency-based Random Perturbation method (Unimatch) exhibited inferior performance compared to the Supervised Baseline (SupBaseline).
Method DICE(%) $\uparrow$ Jaccard(%) $\uparrow$
SupBaseline 73.74 58.4
Unimatch(baseline model) 68.92 52.57
Ours 74.58 59.46
Table 2: Comparison with baseline on the ISIC-2017 dataset, using 10% labeled data.SupBaseline denotes the performance achieved by supervised leveraging solely labeled imagery within the scope of the study. Our method, AIGCMatch, has the best performance.
In contrast, our approach demonstrated a substantial improvement over Unimatch, achieving notable advancements over the SupBaseline. Specifically, our method outperformed the supervised baseline by 0.84% in terms of the Dice Similarity Coefficient and by 1.06% in the Jaccard index. These results indicate the effectiveness of our method in enhancing segmentation performance on the ISIC-2017 dataset, suggesting its potential for broader applications in medical image analysis.
4.4 Ablation study
@@
4.4.1 Ablation study on Perturbation at the image and feature level
@@
Table 3: Comparison with 4 different methods to add perturbation. Random represents adding perturbation randomly while AttPerb represents the attention-guided perturbation method.
@@
4.4.2 Ablation study on Perturbation methods
@@
Table 4: Comparison with 4 different methodologies on the ACDC dataset, using 1, 3, and 7 labeled cases.
@@
5 Conclusions
In this study, we introduced an easy yet efficient semi-supervised segmentation framework termed Attention-Guided weak-to-strong Consistency Match (AIGCMatch), aiming at enhancing the performance of medical image segmentation. By employing attention-guided perturbation strategies at both the image and feature levels, AIGCMatch achieves weak-to-strong consistency and preserves the structural information of medical images as well. Experiments on the ACDC dataset and ISIC-2017 dataset demonstrate the effectiveness of our approach, and we achieve the performance of SOTA on the ACDC dataset. Our work demonstrates the efficacy of attention-guided perturbation strategies in medical image segmentation, improving model robustness by introducing structured and semantically relevant perturbations within the model’s decision-making process.
However, there is still room for further improvement and expansion. Future work will explore the application of AIGCMatch to a broader range of medical image datasets and investigate its performance on more complex and diverse medical images.

### v3

@@ IV-A2 Implementation Details
To ensure a fair comparison with prior work, we just adopted a simple U-Net architecture as the backbone for our segmentation model. Following previous work [15, 49], we employ the Dice Similarity Coefficient, the Jaccard Index, the 95th percentile Hausdorff Distance (95HD), and the Average Surface Distance (ASD) to evaluate the segmentation accuracy of our medical image analysis model. We initialized the learning rate at 0.01 and used an SGD optimizer. The model was trained for 300 epochs under a learning rate scheduler. The original images were resized to vary between 0.5 and 2.0 times their original dimensions, and subjected to cropping and flipping to obtain their weakly perturbed versions. The weights for the loss function, denoted as $\lambda$ and $\mu$, are set to 0.5.
IV-B Semi-supervised image segmentation in ACDC
@@
In the ACDC dataset, we compared our experimental results with those of multiple models, consistently achieving superior results. In comparison to the previous state-of-the-art (SOTA), our method demonstrated improvements across all four metrics.
@@
It should be noted that, compared to previous methodologies, our approach has yielded outcomes equivalent or superior to those of other models that have used 10% labeled data, although it only employs 5% labeled data. Furthermore, compared to previous methods, DiffRect, which labeled 5% labeled data, our approach has achieved comparable outcomes with only 1% of the data, which has often been overlooked due to its poor performance in previous studies. This substantiates the superior performance and substantial potential of AIGCMatch in the domain of medical image segmentation.
IV-C Ablation study
@@
IV-C1 Ablation study on Perturbation at the image and feature level
@@
TABLE II: Comparison with 4 different methods to add perturbation. Random represents adding perturbation randomly while AttPerb represents the attention-guided perturbation method.
@@
IV-C2 Ablation study on Perturbation methods
@@
TABLE III: Comparison with 4 different methodologies on the ACDC dataset, using 1, 3, and 7 labeled cases.
@@
V Conclusions
In this study, we introduced an easy yet efficient semi-supervised segmentation framework termed Attention-Guided weak-to-strong Consistency Match (AIGCMatch), aiming at enhancing the performance of medical image segmentation. By employing attention-guided perturbation strategies at both the image and feature levels, AIGCMatch achieves weak-to-strong consistency and preserves the structural information of medical images as well. Experiments on the ACDC dataset demonstrate the effectiveness of our approach, and we achieve the performance of SOTA on the ACDC dataset.
Future work will explore the application of AIGCMatch to a broader range of medical image datasets and investigate its performance on more complex and diverse medical images.